# Complete $k$-partite entanglement measure

### v2

@@ Abstract
The $k$-partite entanglement, which focus on at most how many particles in the global system are entangled but separable from other particles, is complementary to the $k$-entanglement that reflects how many splitted subsystems are entangled under partitions of the systems in characterizing multipartite entanglement. Very recently, the theory of the complete $k$-entanglement measure has been established in [Phys. Rev. A 110, 012405 (2024)]. Here we investigate whether we can define the complete measure of the $k$-partite entanglement. Consequently, with the same spirit as that of the complete $k$-entanglement measure, we present the axiomatic conditions that a complete $k$-partite entanglement measure should require. Furthermore, we present two classes of $k$-partite entanglement measures and discuss the completeness of them as illustrations.
preprint: APS/123-QED
I Introduction
@@
In 2005, Gühne et al. introduced the $k$-partite entanglement in Ref. [1]. It is closely related to the $k$-producible state: if a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. While the $k$-entanglement reflects how many splitted subsystems are entangled under partitions of the systems, the $k$-partite entanglement concentrate on at most how many particles in the global system are entangled but separable from other particles. It has been shown that $k$-producibility plays a crucial role in both quantum nonlocality [2, 3, 4, 5] and quantum metrology [6]. Particularly, $k$-producibly entangled states for larger $k$ exhibit higher sensitivity in phase estimation [7, 8, 9].
@@
Clearly, these two ways of exhibiting entanglement, i.e., the $k$-entanglement and the the $k$-partite entanglement, are complementary to each other in characterizing the multipartite entanglement which remains challenging to understand undeniably since the complexity increases substantially with the number of parties [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31]. Recently, the $k$-partite entanglement measure based on concurrence have been presnted [30, 31]. Very recently, we established the theory of the complete $k$-entanglement measure in Ref. [27]. It was shown that, in the framework of the complete measure of quantum correlation, the distribution of the correlation could be depicted exhaustively since the correlation could be compared not only between the global system and the subsystem (or the systems under arbitrary partition) but also between different subsystems (or the systems under arbitrary partition) [23, 32, 25, 33, 26, 27]. Along this line, the aim of this paper is to discuss how can we define the complete $k$-partite entanglement measure.
@@
The rest of the paper is arranged as follows. We review the concept of $k$-partite entanglement, the $k$-partite entanglement measures proposed in Ref. [30, 31], and the coarsening relation of multipartite partitions in Sec. II. In Sec. III, we present the definition of the complete $k$-partite entanglement measure, and then give two general ways of constructing $k$-partite entanglement measures and discuss whether they are complete in Sec. IV. Sec. V lists some examples of $k$-partite entanglement measures according to the ways in Sec. IV. Finally, in Sec. VI, we summarize the results of the paper.
II Notations and Preliminaries
@@
For convenience of discussing the complete measure of the $k$-partite entanglement in the next sections, we review some basic notations and terminologies in Sec II.1, and introduce the $k$-partite entanglement measure in literature so far in Sec. II.2. We then introduce the coarsening relation of the multipartite partitions which is necessary when we discuss the completeness of a multipartite quantum correlation measure (also see in Ref. [25, 26]).
@@
We fix some notations first. We denote by $A_{1}A_{2}\cdots A_{n}$ an $n$-partite quantum system. Let $X_{1}|X_{2}|\cdots|X_{m}$ be an $m$-partition of $A_{1}A_{2}\cdots A_{n}$ (for instance, partition $AB|C|DE$ is a $3$-partition of the 5-particle system $ABCDE$ with $X_{1}=AB$, $X_{2}=C$ and $X_{3}=DE$. The case of $m=n$ is just the trivial case that without any partition. So $m<n$ in general unless otherwise specified). We denote by $\varDelta(X_{t})$ the number of subsystems contained in $A_{t}$, for instance, for the 3-partition $AB|C|DE$ of $ABCDE$, $\varDelta(X_{1})=\varDelta(AB)=2$, $\varDelta(X_{2})=\varDelta(C)=1$ and $\varDelta(X_{3})=\varDelta(DE)=2$. If $\varDelta(X_{t})\leqslant k$ for any $1\leqslant t\leqslant m$, we call it a $k$-fineness partition. We denote by $\Gamma_{\!k}^{f}$ the set of all $k$-fineness partitions of the given system $A_{1}A_{2}\cdots A_{n}$.
II.1 $k$-partite entanglement
@@
A pure state $|\psi\rangle$ of an $n$-partite system $A_{1}A_{2}\cdots A_{n}$ with state space $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ is called $k$-producible ($11\leqslant k\leqslant n-$ 1), if it can be represented as [1]
@@
under some $k$-fineness partition $X_{1}|X_{2}|\cdots|X_{m}$ of $A_{1}A_{2}\cdots A_{n}$. Let $\mathcal{S}^{X}$ be the set of all density operators acting on the state space $\mathcal{H}^{X}$. For mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, if it can be written as a convex combination of $k$-producible pure states, i.e., $\rho=\sum\limits_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangle$s are $k$-producible, it is called $k$-producible, where the pure state $|\psi_{i}\rangle$s might be $k$-producible in different $k$-fineness partitions. If a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. Note that if $|\psi\rangle$ admits the form as in Eq. (1), it is called $m$-separable. $|\psi\rangle$ is $m$-entangled if it is not $m$-separable. An $n$-partite mixed state $\rho$ is $m$-separable if it can be written as $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangle$s are $m$-separable, wherein the contained $\{|\psi_{i}\rangle\}$ can be $m$-separable with respect to different $m$-partitions. Otherwise, it is called $m$-entangled. By definition, the $k$-partite entanglement is different from the $k$-entanglement in general, but they are equivalent only in some special cases. For example, the $n$-partite entangled state is just the genuine multipartite entangled state and the one-producible states coincide with the fully separable states. If $|\psi\rangle^{ABC}$ is a genuine entangled state, then $|\psi\rangle^{ABC}|\psi\rangle^{D}|\psi\rangle^{E}|\psi\rangle^{F}$ is four-separable and three-partite entangled state. Also note that, a state of which some reduced state of $m$ parties is genuinely entangled, contains $m$-partite entanglement, but not vice versa in general [1]. For more clarity, we compare $3$-partite entangled pure state with $3$-entangled pure state in Fig. 1.
@@
Figure 1: (color online). (a) $3$-partite entangled pure state $|\Psi\rangle=|\psi\rangle^{A_{1}A_{2}A_{3}}|\psi\rangle^{B_{1}B_{2}\cdots B_{p}}\otimes\cdots\otimes|\psi\rangle^{X_{1}X_{2}\cdots X_{q}}|\phi\rangle^{A_{1}A_{2}}|\phi\rangle^{B_{1}B_{2}}\otimes\cdots\otimes|\phi\rangle^{X_{1}X_{2}}|\varphi\rangle^{A}|\varphi\rangle^{B}\cdots|\varphi\rangle^{X}$, where $|\psi\rangle^{A_{1}A_{2}A_{3}}$, $|\psi\rangle^{B_{1}B_{2}\cdots B_{p}}$, $\dots$, $|\psi\rangle^{X_{1}X_{2}\cdots X_{q}}$ are genuinely entangled states, $3\leqslant p\leqslant q$, $|\phi\rangle^{A_{1}A_{2}}$, $|\phi\rangle^{B_{1}B_{2}}$, $\dots$, $|\phi\rangle^{X_{1}X_{2}}$ are entangled states. In fact, if one of $|\psi\rangle^{A_{1}A_{2}A_{3}}$, $|\psi\rangle^{B_{1}B_{2}\cdots B_{p}}$, $\dots$, $|\psi\rangle^{X_{1}X_{2}\cdots X_{q}}$ is genuinely entangled, $|\Psi\rangle$ is also $3$-partite entangled. Here we just take the general form of a $3$-partite entangled pure state. (b) $|\Phi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ with $k$, $l\geqslant 0$, $k+l\geqslant 3$, is a $3$-entangled pure state if one of the following is true: (i) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are genuinely entangled states, $k$, $l\geqslant 3$, (ii) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are entangled states, $k=l=2$, (iii) If $k=0$ or $l=0$, $|\Phi\rangle$ is genuinely entangled.
@@
A pure state $|\phi\rangle$ is said to be genuinely $k$-producible (or genuinely $k$-partite entangled) [1] if it is $k$-producible but not $(k-\!1)$-producible. A mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is genuinely $k$-producible if it is $k$-producible and for any $k$-producible pure states ensemble of $\rho$, $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, there is at least one $|\psi_{i}\rangle$ is genuinely $k$-producible.
@@
II.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
A positive function $E_{(k)}:\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}\longrightarrow\mathbb{R}_{+}$ is called a $k$-partite entanglement measure ($k$-PEM) if it fulfills: (i) $E_{(k)}(\rho)=0$ for any $\rho\in\mathcal{S}_{P(k-1)}$ and $E_{(k)}(\rho)>0$ for any $\rho\in\mathcal{S}\setminus\mathcal{S}_{P(k-1)}$, (ii) $E_{(k)}(\rho)$ does not increase under $n$-partite local operations and classical communication (LOCC), namely, $E_{(k)}(\varepsilon(\rho))\leq E_{(k)}(\rho)$ for any $n$-partite LOCC $\varepsilon$. Item (ii) guarantees that $E_{(k)}$ is invariant under local unitary operations. In addition, a $k$-PEM $E_{(k)}$ on $\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is convex and non-increasing on average under $n$-partite LOCC, it is call a $k$-partite entanglement monotone ($k$-PEMo).
@@ II.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
where $\rho_{X_{t}}={\rm Tr}_{{\overline{X}}_{t}}(|\psi\rangle\langle\psi|)$, ${\overline{X}}_{t}$ is the complement of subsystem $X_{t}$, the minimum is taken over all the $(k\!-\!1)$-fineness partitions in $\Gamma_{\!k-\!1}^{f}$.
@@
and the $k$-PEMo via the $\alpha$-concurrence was by [31]
@@
where the minimum is taken over all the $(k\!-\!1)$-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. Note here that, the notations here are different from $E_{q-k}$ and $E_{\alpha-k}$ in Ref. [31]: $C_{q(k+1)}=E_{q-k}$, $C_{\alpha(k+1)}=E_{\alpha-k}$. Ref. [31] also gave the following two $k$-PEMos:
$C_{G,q(k)}(|\phi\rangle)=\left(\frac{\prod\limits_{\gamma_{i}\in\Gamma_{\!k-\!1}^{f}}\left[\sum\limits_{t=1}^{m_{i}}(1-{\rm Tr}\rho_{X_{t(i)}}^{q})\right]}{\prod_{i=1}^{\left|\Gamma_{\!k-\!1}^{f}\right|}m_{i}}\right)^{\frac{1}{2\left|\Gamma_{\!k-\!1}^{f}\right|}},~{}~{}~{}$ (6)
@@
where $\rho_{X_{t(i)}}$ is the reduced density operator with respect to subsystem ${X_{t(i)}}$, and $m_{i}$ refers to $\gamma_{i}$ is a $m_{i}$-partition, $\left|\Gamma_{\!k-\!1}^{f}\right|$ is the cardinal number of $\Gamma_{\!k-\!1}^{f}$. The notations in Eqs. (6), (7) are different from $\varepsilon_{q-k}$ and $\varepsilon_{\alpha-k}$ in Ref. [31]: $C_{G,q(k+1)}=\varepsilon_{q-k}$, $C_{G,\alpha(k+1)}=\varepsilon_{\alpha-k}$.
II.3 Coarsening relation of multipartite partitions

### v3

@@ Abstract
The $k$-partite entanglement, which focus on at most how many particles in the global system are entangled but separable from other particles, is complementary to the $k$-entanglement that reflects how many splitted subsystems are entangled under partitions of the systems in characterizing multipartite entanglement. Very recently, the theory of the complete $k$-entanglement measure has been established in [Phys. Rev. A 110, 012405 (2024)]. Here we investigate whether we can define the complete measure of the $k$-partite entanglement. Consequently, with the same spirit as that of the complete $k$-entanglement measure, we present the axiomatic postulates that a complete $k$-partite entanglement measure should require. Furthermore, we present two classes of $k$-partite entanglement measures and show that one is complete while the other one is unified but not complete except for the case of $k=2$.
Keywords: $k$-partite entanglement, $k$-partite entanglement measure, Unified/Complete $k$-partite entanglement measure
1 Introduction
@@
In 2005, Gühne et al. introduced the $k$-partite entanglement in Ref. [1]. It is closely related to the $k$-producible state: if a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. While the $k$-entanglement reflects how many splitted subsystems are entangled under partitions of the systems [2, 3], the $k$-partite entanglement concentrate on at most how many particles in the global system are entangled but separable from other particles [1, 4, 5]. It has been shown that $k$-producibility plays a crucial role in both quantum nonlocality [6, 7, 8, 9] and quantum metrology [10]. Particularly, $k$-producibly entangled states for larger $k$ exhibit higher sensitivity in phase estimation [11, 12, 13].
@@
Clearly, these two ways of exhibiting entanglement, i.e., the $k$-entanglement and the $k$-partite entanglement, are complementary to each other in characterizing the multipartite entanglement which remains challenging to understand undeniably since the complexity increases substantially with the number of parties [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Recently, the $k$-partite entanglement measure based on concurrence have been presnted [4, 5]. Very recently, we established the theory of the complete $k$-entanglement measure in Ref. [2]. It was shown that, in the framework of the complete measure of quantum correlation, the distribution of the correlation could be depicted exhaustively since the correlation could be compared not only between the global system and the subsystem (or the systems under arbitrary partition) but also between different subsystems (or the systems under arbitrary partition) [2, 21, 23, 24, 25, 26]. Along this line, the aim of this paper is to discuss how can we establish the axiomatic postulates for the complete $k$-partite entanglement measure.
@@
The rest of the paper is arranged as follows. We review the concept of $k$-partite entanglement, the $k$-partite entanglement measures proposed in Ref. [4, 5], and the coarsening relation of multipartite partitions in Sec. 2. In Sec. 3, we present the definition of the complete $k$-partite entanglement measure, and then give two general ways of constructing $k$-partite entanglement measures and discuss whether they are complete in Sec. 4. Sec. 5 lists some examples of $k$-partite entanglement measures according to the methods in Sec. 4. Finally, in Sec. 6, we summarize the results of the paper.
2 Notations and Preliminaries
@@
For convenience of discussing the complete measure of the $k$-partite entanglement in the next sections, we review some basic notations and terminologies in Sec 2.1, and introduce the $k$-partite entanglement measures proposed in literature so far in Sec. 2.2. We then introduce the coarsening relation of the multipartite partitions which is necessary when we discuss the completeness of a multipartite quantum correlation measure (also see in Ref. [23, 24]).
@@
We fix some notations first. We denote by $A_{1}A_{2}\cdots A_{n}$ an $n$-partite quantum system. Let $X_{1}|X_{2}|\cdots|X_{m}$ be an $m$-partition of $A_{1}A_{2}\cdots A_{n}$ (for instance, partition $AB|C|DE$ is a $3$-partition of the 5-particle system $ABCDE$ with $X_{1}=AB$, $X_{2}=C$ and $X_{3}=DE$. The case of $m=n$ is just the original $n$-particle system without any other partition, namely, $A_{1}A_{2}\cdots A_{n}$ means $A_{1}|A_{2}|\cdots|A_{n}$. So $m<n$ in general unless otherwise specified). We denote by $\varDelta(X_{t})$ the number of subsystems contained in $X_{t}$, for instance, for the 3-partition $AB|C|DE$ of $ABCDE$, $\varDelta(X_{1})=\varDelta(AB)=2$, $\varDelta(X_{2})=\varDelta(C)=1$ and $\varDelta(X_{3})=\varDelta(DE)=2$. If $\varDelta(X_{t})\leqslant k$ for any $1\leqslant t\leqslant m$, we call it a $k$-fineness partition. We denote by $\Gamma_{\!k}^{f}$ the set of all $k$-fineness partitions of the given system $A_{1}A_{2}\cdots A_{n}$.
2.1 $k$-partite entanglement
@@
A pure state $|\psi\rangle$ of an $n$-partite system $A_{1}A_{2}\cdots A_{n}$ with state space $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ is called $k$-producible ($1\leqslant k\leqslant n-1$), if it can be represented as [1]
@@
under some $k$-fineness partition $X_{1}|X_{2}|\cdots|X_{m}$ of $A_{1}A_{2}\cdots A_{n}$. Let $\mathcal{S}^{X}$ be the set of all density operators acting on the state space $\mathcal{H}^{X}$. For mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, if it can be written as a convex combination of $k$-producible pure states, i.e., $\rho=\sum\limits_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $k$-producible, it is called $k$-producible, where the pure state $|\psi_{i}\rangles$ might be $k$-producible in different $k$-fineness partitions. If a quantum state is not $k$-producible, it is termed $(k+1)$-partite entangled. Note that if $|\psi\rangle$ admits the form as in Eq. (1), it is called $m$-separable. $|\psi\rangle$ is $m$-entangled if it is not $m$-separable. An $n$-partite mixed state $\rho$ is $m$-separable if it can be written as $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$ with $|\psi_{i}\rangles$ are $m$-separable, wherein the contained $\{|\psi_{i}\rangle\}$ can be $m$-separable with respect to different $m$-partitions. Otherwise, it is called $m$-entangled. By definition, the $k$-partite entanglement is different from the $k$-entanglement in general, but they are equivalent only in some special cases. For example, the $n$-partite entangled state is just the genuine multipartite entangled state and the one-producible state coincides with the fully separable state. If $|\psi\rangle^{ABC}$ is a genuine entangled state, then $|\psi\rangle^{ABC}|\psi\rangle^{D}|\psi\rangle^{E}|\psi\rangle^{F}$ is four-separable and three-partite entangled state. Also note that, a state of which some reduced state of $m$ parties is genuinely entangled, contains $m$-partite entanglement, but not vice versa in general [1]. For more clarity, we compare $3$-partite entangled pure state with $3$-entangled pure state in Fig. 1.
@@
Figure 1: (color online). (a) 33-partite entangled pure state $|\Psi\rangle=|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}\cdots|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}|\phi\rangle^{A_{1}A_{2}}|\phi\rangle^{B_{1}B_{2}}\cdots|\phi\rangle^{X_{1}X_{2}}|\varphi\rangle^{A}|\varphi\rangle^{B}\cdots$ $|\varphi\rangle^{X}$, where $|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}$, $|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}$, $\dots$, $|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}$ are genuinely entangled states, $3\leqslant p\leqslant q$, $|\phi\rangle^{A_{1}A_{2}}$, $|\phi\rangle^{B_{1}B_{2}}$, $\dots$, $|\phi\rangle^{X_{1}X_{2}}$ are entangled states. In fact, if one of $|\psi\rangle^{A^{\prime}_{1}A^{\prime}_{2}A^{\prime}_{3}}$, $|\psi\rangle^{B^{\prime}_{1}B^{\prime}_{2}\cdots B^{\prime}_{p}}$, $\dots$, $|\psi\rangle^{X^{\prime}_{1}X^{\prime}_{2}\cdots X^{\prime}_{q}}$ is genuinely entangled, $|\Psi\rangle$ is also 33-partite entangled. Here we just take the general form of a $3$-partite entangled pure state. (b) $|\Phi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ with $k$, $l\geqslant 0$, $k+l\geqslant 3$, is a 33-entangled pure state if one of the following is true: (i) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are genuinely entangled states, $k$, $l\geqslant 3$, (ii) $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}$ and $|\psi\rangle^{B_{1}B_{2}\cdots B_{l}}$ are entangled states, $k=l=2$, (iii) If $k=0$ or $l=0$, $|\Phi\rangle$ is genuinely entangled.
@@
A pure state $|\phi\rangle$ is said to be genuinely $k$-producible (or genuinely $k$-partite entangled) [1] if it is $k$-producible but not (k1)(k-\!1)-producible. A mixed state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is genuinely $k$-producible if it is $k$-producible and for any $k$-producible pure states ensemble of $\rho$, $\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|$, there is at least one $|\psi_{i}\rangle$ is genuinely $k$-producible.
@@
2.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
A positive function $E_{(k)}:\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}\longrightarrow\mathbb{R}_{+}$ is called a $k$-partite entanglement measure ($k$-PEM) if it fulfills: (i) $E_{(k)}(\rho)=0$ for any $\rho\in\mathcal{S}_{P(k-1)}$ and $E_{(k)}(\rho)>0$ for any $\rho\in\mathcal{S}\setminus\mathcal{S}_{P(k-1)}$, (ii) $E_{(k)}(\rho)$ does not increase under $n$-partite local operations and classical communication (LOCC), namely, $E_{(k)}(\varepsilon(\rho))\leq E_{(k)}(\rho)$ for any $n$-partite LOCC $\varepsilon$. Item (ii) guarantees that $E_{(k)}$ is invariant under local unitary operation. In addition, a $k$-PEM $E_{(k)}$ on $\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ is convex and non-increasing on average under $n$-partite LOCC, it is call a $k$-partite entanglement monotone ($k$-PEMo). A $k$-PEM/$k$-PEMo $E_{(k)}$ is called a genuine $k$-PEM/$k$-PEMo if $E_{(k)}(\rho)>0$ but $E_{(k+1)}(\rho)=0$ for any genuinely $k$-partite entangled state $\rho$.
@@ 2.2 $k$-partite entanglement measures via $q$-concurrence and $\alpha$-concurrence
where $\rho_{X_{t}}={\rm Tr}_{{\overline{X}}_{t}}(|\psi\rangle\langle\psi|)$, ${\overline{X}}_{t}$ is the complement of subsystem $X_{t}$, the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. For mixed states, it is defined by the convex-roof extension, i.e.,
$C_{(k)}(\rho)=\min\limits_{p_{i},|\psi_{i}\rangle}\sum_{i}p_{i}C_{(k)}(|\psi_{i}\rangle),$
where the minimum runs over all ensembles $\{p_{i},|\psi_{i}\rangle|\rho=\sum_{i}p_{i}|\psi_{i}\rangle\langle\psi_{i}|\}$. In what follows, we give only the measures for pure states, for the case of mixed states they are all defined by the convex-roof extension with no further statement. Obviously, any measure that is defined in this way is convex.
@@
and the $k$-PEMo via the $\alpha$-concurrence was expressed by [5]
@@
where the minimum is taken over all the (k1)(k\!-\!1)-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. Note here that, the notations here are different from $E_{q-k}$ and $E_{\alpha-k}$ in Ref. [5] ($C_{q(k+1)}=E_{q-k}$, $C_{\alpha(k+1)}=E_{\alpha-k}$). Ref. [5] also gave the following two $k$-PEMos:
$C_{G,q(k)}(|\phi\rangle)=\left(\frac{\prod\limits_{\gamma_{i}\in\Gamma_{\!k-\!1}^{f}}\left[\sum\limits_{t=1}^{m_{i}}(1-{\rm Tr}\rho_{X_{t(i)}}^{q})\right]}{\prod_{i=1}^{\left|\Gamma_{\!k-\!1}^{f}\right|}m_{i}}\right)^{\frac{1}{2\left|\Gamma_{\!k-\!1}^{f}\right|}}$ (6)
@@
where $\rho_{X_{t(i)}}$ is the reduced density operator with respect to subsystem $X_{t(i)}$}, and $m_{i}$ refers to $\gamma_{i}$ is a $m_{i}$-partition, $\left|\Gamma_{\!k-\!1}^{f}\right|$ is the cardinal number of $\Gamma_{\!k-\!1}^{f}$. The notations in Eqs. (6), (7) are different from $\varepsilon_{q-k}$ and $\varepsilon_{\alpha-k}$ in Ref. [5] ($C_{G,q(k+1)}=\varepsilon_{q-k}$, $C_{G,\alpha(k+1)}=\varepsilon_{\alpha-k}$).
2.3 Coarsening relation of multipartite partitions